OpenStudy (anonymous):

Choose the equation of a line with a slope of 3/4 and a y-intercept of (0, 1).

OpenStudy (anonymous):

\[ y=mx+b \]So \(m=3/4\) and \(b=1\).

OpenStudy (anonymous):

is there more? these are the answers A.4x - 3y = - 1 B.3x + 4y = 1 C.3x - 4y = - 4 D.None of these. E.3x - 4y = -2

OpenStudy (anonymous):

its looks like it wants you to put it into standard form, which is: \[Ax+Bx=C\]so we have \[y=3/4x+1\] now we just have to move our variables around to match standard form

OpenStudy (anonymous):

Well you have to change the equation a bit. We start out as \[ y=\frac 34x +1 \]

OpenStudy (anonymous):

\[Ax + By = C\]

OpenStudy (anonymous):

Manipulate it with subtraction and multiplication.

OpenStudy (anonymous):

@k26 Do you know how to get \(3/4x\) onto the other side of the equation?

OpenStudy (anonymous):

yes got it thanks

OpenStudy (anonymous):

Solve the following system of equations for x: -x + 9y = - 4 5x - 45y = 2